OpenStudy (anonymous):

Solve for x: (x + 20)/2 =3x ^ x + 20 over 2

OpenStudy (texaschic101):

(x + 20) / 2 = 3x --- multiply both sides by 2 x + 20 = 6x -- subtract x from both sides 20 = 6x - x -- combine like terms 20 = 5x -- divide both sides by 5 4 = x check.. (x + 20) / 2 = 3x (4 + 20) / 2 = 3(4) 24/2 = 12 12 = 12 (correct) x = 4 any questions ?